OpenStudy (anonymous):

What is the simplified form of each expression? 7x^–8 × 6x^3 a. 42/x^5 b. 1/42x^5 c. 42x^11 d. 13x^–5 (–2x8) · 3y9 · 2x4 a. 3x^12y^9 b. –12x^72y^9 c. –12xy^21 d. –12x^y^9 Find the simplified form of the expression. Give your answer in scientific notation. (8 x 10^7) (7 x 10^4) a. 1.5 × 10^12 b. 5.6 × 10^12 c. 1.5 x 10^29 d. 5.6 x 10^29 (7 × 10^–4)(9 × 10^–10) a. 6.3 × 10^–13 b. 63 × 10^–13 c. 6.3 × 10^–15 d. 6.3 × 10^–16

OpenStudy (phi):

it is not good to post multiple questions

OpenStudy (anonymous):

Can you do the first one?

OpenStudy (anonymous):

first one --- a.

OpenStudy (phi):

first, the numbers can be multiplied . 6*7=42 so you can rule out some of the choices that way

OpenStudy (anonymous):

Thanks GiGlgo3

OpenStudy (phi):

second, there is the idea that anything divided by itself is 1 so if you see x/x \[ \frac{x}{x} =1 \] or \[ \frac{x\cdot x}{x\cdot x } =1 \] 2nd, remember that x^3 (for example) means x*x*x (x times itself 3 times) I would rewrite the x's by flipping the negative one to get \[ \frac{x^3}{x^8} \] if we expand it we get \[ \frac{x*x*x}{x*x*x*x*x*x*x*x} \] now for every x on the top and bottom we can divide (because x/x=1) \[ \cancel{\frac{x*x*x}{x*x*x}}\frac{1}{x*x*x*x*x} \] so the x's are (using the short hand \[ \frac{1}{x^5} \] of course we still have the 6*7=42 so the answer is \[ \frac{42}{x^5} \]

OpenStudy (anonymous):

Wow! Thank you!

OpenStudy (phi):

That is how it works, but people pretty quickly figure out that if they combine the exponents -8+3 = -5 you get the answer right away

OpenStudy (phi):

so the answer could be written 42 x^(-5) or 42/x^5

OpenStudy (anonymous):

Can you help me with the 3rd and 4th one?

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

3. ---- b.

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry if its trouble...I just like to know how you got it:) It helps me understand:)

OpenStudy (anonymous):

im here, im using notepad. (:

OpenStudy (anonymous):

Okay:)

OpenStudy (anonymous):

Can you do that last question...and explain?

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

first step 10 to the seventh power is 10000000 times 8 is 80000000 10 to the fourth power is 10000 times 7 is 70000 80000000 times 70000 is 5600000000000

OpenStudy (anonymous):

Wow your good

OpenStudy (anonymous):

step 2: 10 to the 12th power is 1000000000000 (hint 12 zeros) 1000000000000 times 1.5 is 500000000000... thats not right now is it?

OpenStudy (anonymous):

step 3: 10 to the 12th power is 1000000000000 (hint 12 zeros) 1000000000000 times 5.6 is 5600000000000.. that's right! so b. is the right answer... (:

OpenStudy (anonymous):

can you try and do five, or do you need a walkthrough again??

OpenStudy (anonymous):

So the last problems first step would be 630000000000000000?

OpenStudy (anonymous):

let me think

OpenStudy (anonymous):

The zeros get me confused! lol

OpenStudy (anonymous):

yea, you are riight

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Okay...I need help on the last steps though

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

Can you help me with the last steps?

OpenStudy (anonymous):

So its 6.3 x 10^-13?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I cant believe i got that!

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

bye, i have to go, remember, just message me. (:

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

Oh I meant thank you before you said message me. Okay byeee!